Jasper Company has sales on account and for cash. Specifically, 59% of its sales are on account and 41% are for cash. Credit sales are collected in full in the month following the sale. The company forecasts sales of $531,000 for April, $541,000 for May, and $566,000 for June. The beginning balance of Accounts Receivable is $292,400 on April 1 Prepare a schedule of budgeted cash receipts for April, May, and June April May June Cash sales 41%, $ 217,7|of $ 221,810r $ 232,060 333,940 $ 531,000$ 541,000 $ 566,000 Sales on account 59% 313,290 319,190 Total sales JASPER COMPANY Schedule of Cash Receipts For April, May, and June April May June Cash receipts from Cash sales $ 217,710$ 221,810 $ 232,060 319,190 $ 510,110$ 535,100 $ 551,250 Collection of accounts receivable 292,400 313,290 Total budgeted cash receipts
Answer #1
April May June
Cash sales 41% 217710 221810 232060
Sales on account 59% 313290 319190 333940
Total sales 531000 541000 566000
Schedule of Cash Receipts
For April, May, and June
April May June
Cash receipts from:
Cash sales 217710 221810 232060
Collections of accounts Receivable 292400 313290 319190
Total budgeted cash receipts 510110 535100 551250
Formulas used:
April May June
Cash sales =531000*41% =541000*41% =566000*41%
Sales on account =531000*59% =541000*59% =566000*59%
=217710+313290 =221810+319190 =232060+333940